In a radio wave if E=10^-4 V/m, Then find the magnitude of poynting vector?
1
Expert's answer
2020-06-15T10:27:51-0400

Since, in a radio wave EB\overrightarrow{E}\bot \overrightarrow{B} and Poynting vector is defined by

S=1μ0(E×B)    S=1μ0EB\overrightarrow{S}=\frac{1}{\mu_0}(\overrightarrow{E}\times\overrightarrow{B})\\ \implies ||\overrightarrow{S}||=\frac{1}{\mu_0}||\overrightarrow{E}||\cdot||\overrightarrow{B}||

But, we also know that,

B=Ec||\overrightarrow{B}||=\frac{||\overrightarrow{E}||}{c}

Thus, we get

S=1cμ0E2||\overrightarrow{S}||=\frac{1}{c\mu_0}||\overrightarrow{E}||^2

Given, E=E=104V/m||\overrightarrow{E}||=E=10^{-4}V/m

Hence, on plugin the value we get

S=2.655×1011W/m2||\overrightarrow{S}||=2.655\times 10^{-11}W/m^2